An insurance policy requires premium payments of $15 at the beginning of each month for 20 years. Determine the discounted and accumulated values of these payments at j4 = 6%.
Answer #1

Discounted Value = (1+r) AnnualCash Flow * -

where r is the rate of interestfor a compounding period i.e. 6% /12 = 0.5%  

n is the no of compounding periods i.e. 20 years * 12 = 240

= C000 * onz 900 0+1) - T

Discounted Value = $ 2093.71

Discounted Value of Annuity due = $ 2093.71157526 ( 1+ 0.005) = $ 2,104.18

Accumulated Value = (1 + r) - 1 (1 + r) * Periodic Payment *

=(1+0.005)151 + 0.005)240_1 0.005

= $ 6,965.27

NOTE: The answer to your question has been given below/above. If there is any query regarding the answer, please ask in the comment section. If you find the answer helpful, do upvote. Help us help you.